# Explicit Pfaffian Formula for Amplitudes of Fermionic Gaussian Pure States in Arbitrary Pauli Bases

### v1

@@ Abstract
The explicit computation of amplitudes for fermionic Gaussian pure states in arbitrary Pauli bases is a long-standing challenge in quantum many-body physics, with significant implications for quantum tomography, experimental studies, and quantum dynamics. These calculations are essential for analyzing complex properties beyond traditional measures, such as formation probabilities, global entanglement, and entropy in non-standard bases, where exact and computationally efficient methods remain underdeveloped. In particular, having explicit formulas is crucial for optimizing negative log-likelihood functions in quantum tomography, a key task in the NISQ era. In this work, we present an explicit Pfaffian formula (Theorem 1) for determining these amplitudes in arbitrary Pauli bases, utilizing a Pfaffian of a matrix based on qubit parity. Additionally, we introduce a recursive relation (Theorem 2) that links the amplitudes of systems with different qubit counts, enabling scalable calculations for large systems. Together, these results provide a versatile framework for applications in global entanglement, Shannon-Rényi entropy, formation probabilities, and quantum tomography, significantly expanding the computational toolkit for analyzing complex quantum systems.
@@ I Introduction
Fermionic Gaussian states are among the earliest and most widely studied many-body fermionic systems, with applications spanning a range of fields. They serve as exact ground states for coupled spin chains LIEB1961407 , provide approximate ground states for interacting fermions Thouless1960 ; szabo2012modern ; EA2007 ; Kraus2010 ; Terhal2023 , and play an important role in impurity problems Bravyi2017 . Beyond these, they are valuable in quantum simulation, benchmarking quantum computation Knill2001 ; Terhal2002 ; Brod2016 , and modeling the time evolution of free fermionic systems after a quantum quench CEF2011 . While several properties of these states are well-understood—often traceable back to Wick’s theorem, such as entanglement entropy Peschel2001 ; Kitaev2003 and fidelity Zanardi2008 —other characteristics, like efficient tomography Gluza2018 ; Gorman2022 ; Leone2024 , trace distance RZ2023 ; Wang2024 ; Leone2024 , and Shannon-Rényi entropy NRR2024 , have sparked intense recent research. It is evident, however, that calculating all relevant quantities for Gaussian states does not always simplify to Wick’s theorem alone, leading some aspects to be categorized as challenging problems.
@@
Another potential application lies in calculating post-measurement entanglement entropy R2015 and measurement-induced entanglement Potter2024 , which refers to the entanglement entropy of a system following a projective measurement. For post-measurement entanglement entropy, a projective measurement is performed in a local basis on part of the system, followed by a calculation of the bipartite entanglement entropy of the remaining system. In the case of measurement-induced entanglement, one averages overall possible entanglement values following projective measurement. Having explicit formulas for probability amplitudes in alternative bases could enhance numerical algorithms for computing these quantities in free fermionic systems, such as the Ising and XY chains.
In the context of match gates, which are equivalent to free fermions Terhal2002 , it is already established that probabilities can be efficiently simulated classically Brod2016 . This indicates the possible existence of a closed-form expression for the amplitudes. However, existing methods do not appear to provide a straightforward approach to derive an explicit formula for the amplitudes. In this paper, we present an explicit formula for the amplitudes of fermionic Gaussian pure states in an arbitrary local Pauli basis. This formula is expressed in terms of the Pfaffian of a matrix with dimensions $L$ or $L+1$, depending on whether the system has an even or odd number of qubits, respectively, enabling efficient computation of individual amplitudes even for large systems ($L>1000$). We then introduce a recursive formula that relates the amplitudes of an $L$-qubit system to those of systems with $L-2$ and $2$ qubits. In non-computational bases, this relationship is not necessarily within the same basis; for example, the amplitude configurations in the $x$-basis can be related to those in the $y$-basis.
@@
We start the section II with an overview of the properties of free fermion Gaussian pure states, formulated in a way that facilitates the derivation of our results. In section III we proceed by performing a basis transformation to express the amplitudes as an exponential sum. This sum is then simplified to yield a single explicit Pfaffian formula. In section IV, we write an explicit formula for probabilities which can be used for many applications. Finally we conclude the paper in section V with mentioning some possible future generalizations and applications.
@@ IV Formation Probabilities
A simple application of the theorem 1 leads to the following result for the probabilities:
@@
Obviously, the angle $\alpha$ does not play any role in the probability formula. The above formula can be directly applied to numerically or analytically compute formation probabilities, Shannon-Rényi entropies, geometric entropy, transition probabilities in quantum quenches, and to optimize the Negative Log-Likelihood (NLL) in quantum tomographyZR2025 .
V Conclusion
@@
In this work, we derived an explicit Pfaffian formula for computing the amplitudes of fermionic Gaussian pure states in any Pauli basis, along with a recursive relation that connects amplitudes across different qubit sizes. Together, these results establish a robust computational framework that will be valuable for addressing complex quantum information problems, including formation probabilities, global entanglement, and quantum tomography, particularly within the NISQ era. Applying these formulas to specific quantum systems is a natural next step, expected to advance both theoretical and experimental research in Gaussian state analysis. Open questions remain, especially regarding extensions to Gaussian mixed states, where finding similar explicit formulas would broaden the applicability of Gaussian states in many-body physics and quantum computation. The methods introduced here offer essential tools for enhanced state analysis and practical applications in these fields.
@@
Acknowledgements. We thank D. Brod for discussion. MAR thanks CNPq and FAPERJ (Grant No. E-26/210.062/2023) for partial support.

### v3

@@ Abstract
The explicit computation of amplitudes for fermionic Gaussian pure states in arbitrary Pauli bases is a long-standing challenge in quantum many-body physics, with significant implications for quantum tomography, experimental studies, and quantum dynamics. These calculations are essential for analyzing complex properties beyond traditional measures, such as formation probabilities, global entanglement, and entropy in non-standard bases, where exact and computationally efficient methods remain underdeveloped. In addition to these physical applications, having explicit formulas is crucial for optimizing negative log-likelihood functions in quantum tomography, a key task in the NISQ era. In this work, we present an explicit Pfaffian formula (Theorem 1) for determining these amplitudes in arbitrary Pauli bases, utilizing a matrix whose structure reflects the qubit parity. Additionally, we introduce a recursive relation (Theorem 2) that connects amplitudes for systems with varying qubit numbers, enabling scalable computations for large systems. Together, these results provide a versatile framework for studying global entanglement, Shannon-Rényi entropies, formation probabilities, and performing efficient quantum tomography, thereby significantly expanding the computational toolkit for analyzing complex quantum systems. Finally, we utilize our formalism to determine the post-measurement entanglement entropy, reflecting how local measurements alter entanglement, and compare the outcomes with conformal field theory predictions.
@@ I Introduction
Fermionic Gaussian states are among the earliest and most widely studied many-body fermionic systems, with applications spanning a range of fields. They serve as exact ground states for coupled spin chains LIEB1961407 , provide approximate ground states for interacting fermions Thouless1960 ; szabo2012modern ; EA2007 ; Kraus2010 ; Terhal2023 , and play an important role in impurity problems Bravyi2017 . Beyond these, they are valuable in quantum simulation, benchmarking quantum computation Knill2001 ; Terhal2002 ; Brod2016 , and modeling the time evolution of free fermionic systems after a quantum quench CEF2011 . While several properties of these states are well-understood—often traceable back to Wick’s theorem, such as entanglement entropy Peschel2001 ; Kitaev2003 and fidelity Zanardi2008 —other characteristics, like efficient tomography Gluza2018 ; Gorman2022 ; Leone2024 , trace distance RZ2023 ; Wang2024 ; Leone2024 , and Shannon-Rényi entropy NRR2024 , have sparked intense recent research. It is evident, however, that calculating all relevant quantities for Gaussian states does not always simplify to Wick’s theorem alone, leading some aspects to be categorized as challenging problems. For an overview of both analytical and numerical techniques, we refer the reader to Ref. Surace2022 , and references therein.
@@
Another potential application lies in calculating post-measurement entanglement entropy R2015 ; Baweja2024 and measurement-induced entanglement Potter2024 , which refers to the entanglement entropy of a system following a projective measurement. For post-measurement entanglement entropy, a projective measurement is performed in a local basis on part of the system, followed by a calculation of the bipartite entanglement entropy of the remaining system. In the case of measurement-induced entanglement, one averages overall possible entanglement values following projective measurement. Having explicit formulas for probability amplitudes in alternative bases could enhance numerical algorithms for computing these quantities in free fermionic systems, such as the Ising and XY chains.
Despite their wide applicability, fermionic Gaussian states constitute a restricted subset of quantum states. They are efficiently classically simulable, much like stabilizer states in spin systems, and typically fail to capture non-Gaussian correlations. These limitations mean that Gaussian states are not suitable for describing certain interacting quantum phases or for achieving computational advantage in the general case. Accordingly, our framework is confined to the domain of free or weakly interacting systems, where Gaussian descriptions remain powerful. In this context, our explicit methods provide valuable tools for state characterization, efficient simulation, and benchmarking. Moreover, the insights and structural results presented here may inform future extensions toward non-Gaussian ansätze or hybrid methods capable of addressing more complex quantum regimes.
Within this setting, it is known that certain classes of fermionic systems—such as those described by match gates—can be simulated efficiently on a classical computer Terhal2002 ; Brod2016 . This suggests that explicit, closed-form expressions for measurement amplitudes should exist, at least in principle. However, existing methods do not appear to provide a straightforward approach to derive an explicit formula for the amplitudes. In this paper, we present an explicit formula for the amplitudes of fermionic Gaussian pure states in an arbitrary local Pauli basis. This formula is expressed in terms of the Pfaffian of a matrix with dimensions $L$ or $L+1$, depending on whether the system has an even or odd number of qubits, respectively, enabling efficient computation of individual amplitudes even for large systems ($L>1000$). We then introduce a recursive formula that relates the amplitudes of an $L$-qubit system to those of systems with $L-2$ and $2$ qubits. In non-computational bases, this relationship is not necessarily within the same basis; for example, the amplitude configurations in the $x$-basis can be related to those in the $y$-basis.
@@
We start the section II with an overview of the properties of free fermion Gaussian pure states, formulated in a way that facilitates the derivation of our results. In section III, we proceed by performing a basis transformation to express the amplitudes as an exponential sum. This sum is then simplified to yield a single explicit Pfaffian formula. In section IV, we write an explicit formula for probabilities that can be used for many applications. Building on this, section V shows how our framework allows for the efficient computation of post-measurement entanglement entropy in arbitrary Pauli bases. Finally, we conclude the paper in section VI by mentioning some possible future generalizations and applications.
@@ IV Formation Probabilities
A simple application of the theorem 1 leads to the following result for the formation probabilities:
@@
Here, $P_{\mathcal{S}}(\boldsymbol{\phi},\boldsymbol{\theta},\boldsymbol{\alpha})$ denotes the probability of obtaining the outcome $\left|\mathcal{S}\right\rangle=\left|s_{1},s_{2},\dots,s_{L}\right\rangle$ after a projective measurement in the basis specified by $(\boldsymbol{\phi},\boldsymbol{\theta},\boldsymbol{\alpha})$. Obviously, the angle $\alpha$ does not play any role in the probability formula. The above formula can be directly applied to numerically or analytically compute formation probabilities, Shannon-Rényi entropies, geometric entropy, transition probabilities in quantum quenches, and to optimize the Negative Log-Likelihood (NLL) in quantum tomographyZR2025 .
[…]
Figure 1: Schematic of a one-dimensional chain with periodic boundary conditions, forming a closed ring. The chain is divided into two regions, $A$ and $B$, where $A=A_{1}\cup A_{2}$ and $B=B_{1}\cup B_{2}$. A projective measurement is applied on the $B_{1}$ and $B_{2}$ subsystems. We then study how the entanglement between $A_{1}$ and $A_{2}$ depends on the distance separating them along the ring.
V Post measurement entanglement entropy
Post-measurement entanglement entropy reveals how local measurements reshape quantum correlations within a system. While CFT provides predictions in specific cases Rajabpour2015 ; Antonini2023 ; Numasawa2016 ; Guo2019 ; Antonini2022 ; Weinstein2023 , calculating this quantity in a general measurement basis is challenging. Our explicit formulas offer a practical and general method to compute the entanglement entropy after projective measurements in arbitrary Pauli bases, making it possible to extract scaling behavior and decay exponents.
To demonstrate this, consider a bipartition of the system into subsystems $A$ and $B$. We perform a local projective measurement on subsystem $B$ in the basis defined by the angles $(\boldsymbol{\phi},\boldsymbol{\theta},\boldsymbol{\alpha})$. A schematic illustration of this setup is shown in Figure 1. In our framework, the full state is expressed in terms of a configuration $\mathcal{S}$ that decomposes as
$\mathcal{S}=(\mathcal{S}_{A},\mathcal{S}_{B}),$
with the probability of the full configuration given by Eq. (11). Consequently, the probability of obtaining a specific measurement outcome $\mathcal{S}_{B}$ in subsystem $B$ is obtained by summing over all configurations in subsystem $A$:
$P_{\mathcal{S}_{B}}(\boldsymbol{\phi},\boldsymbol{\theta},\boldsymbol{\alpha})=\sum_{\mathcal{S}_{A}}P_{\mathcal{S}}(\boldsymbol{\phi},\boldsymbol{\theta},\boldsymbol{\alpha})\,.$ (12)
Refer to caption(b)
Figure 2: Transverse-field Ising chain at the critical point. The main plots show the power-law decay of $\mathcal{E}_{A_{1}A_{2}}^{(\alpha)}$ as a function of distance $d$, while the insets display the corresponding decay exponent $\eta_{\alpha}$ for different values of $\alpha$. The dashed line represents the extrapolation to the thermodynamic limit ($L\to\infty$). The measurements are in the (a) $\sigma_{z}$ (b) $\sigma_{x}$ basis.
Once the measurement is performed, the overall state collapses to one conditioned on the outcome $\mathcal{S}_{B}$. The post measurement state of subsystem $A$ is then given by
$\left|\psi_{A}^{(\mathcal{S}_{B})}\right\rangle=\frac{1}{\sqrt{P_{\mathcal{S}_{B}}}}\sum_{\mathcal{S}_{A}}a_{(\mathcal{S}_{A},\mathcal{S}_{B})}(\mathbf{R},\boldsymbol{\phi},\boldsymbol{\theta},\boldsymbol{\alpha})\,\left|\mathcal{S}_{A}\right\rangle\,.$ (13)
After the measurement, the wave function collapses into a product state between $A$ and $B$, eliminating entanglement between the two subsystems. However, our interest lies in the residual (or induced) entanglement within subsystem $A$ itself. To capture this, we further partition $A$ into two disjoint parts, $A_{1}$ and $A_{2}$. The reduced density matrix of $A_{1}$ is obtained by tracing over the degrees of freedom in $A_{2}$:
$\rho_{A_{1}}^{(\mathcal{S}_{B})}=\mathrm{tr}_{A_{2}}\left[\left|\psi_{A}^{(\mathcal{S}_{B})}\right\rangle\left\langle\psi_{A}^{(\mathcal{S}_{B})}\right|\right]\,.$ (14)
Finally, the entanglement entropy between $A_{1}$ and $A_{2}$ for a given measurement outcome $\mathcal{S}_{B}$ can be quantified using the Rényi entanglement entropy, defined as
$\mathcal{E}_{A_{1}}^{(\alpha,\mathcal{S}_{B})}=\frac{1}{1-\alpha}\ln\mathrm{tr}\left[\left(\rho_{A_{1}}^{(\mathcal{S}_{B})}\right)^{\alpha}\right]\,,$ (15)
where $\alpha>0$ and $\alpha\neq 1$. In the limit $\alpha\to 1$, the Rényi entropy reduces to the von Neumann entropy, given by
$\mathcal{E}_{A_{1}}^{(\mathcal{S}_{B})}=-\mathrm{tr}\left[\rho_{A_{1}}^{(\mathcal{S}_{B})}\ln\rho_{A_{1}}^{(\mathcal{S}_{B})}\right]\,.$ (16)
To place our findings in a broader context, we first present two key results derived from conformal field theory (CFT) applied to critical spin chains with short-range interactions. These results describe the scaling of post measurement entanglement entropy and the behavior of the decay exponent $\eta(\alpha)$ as a function of the Rényi index $\alpha$ R2016 :
$\mathcal{E}_{A_{1}}^{(\alpha,\mathcal{S}_{B})}(d)\sim d^{-\eta(\alpha)},$ (17)
where $d$ denotes the distance between the two disjoint subsystems $A_{1}$ and $A_{2}$, and the sizes of these subsystems satisfy $|A_{1}|,|A_{2}|\ll|B|$. The asymptotic exponent is given by
$\eta(\alpha)=\begin{cases}4\alpha\Delta_{1},&\text{if }\alpha<1\\ 4\Delta_{1},&\text{if }\alpha\geq 1\end{cases}$ (18)
Here, $\Delta_{1}$ is the scaling dimension associated with the operator that creates the post-measurement state. Its value is determined by the boundary condition imposed by the configuration $\mathcal{S}_{B}$ from a renormalization group perspective and must be computed explicitly on the discrete spin chain. These expressions hold for general spin chains at criticality, provided the system admits a CFT description. Although the value of $\Delta_{1}$ depends on the choice of measurement basis, the outcome of measurement and the underlying critical model it still shows some level of universality that we will study here. In the critical transverse-field Ising chain, the scaling dimension $\Delta_{1}$ has been numerically studied in the $\sigma^{z}$ basis R2016 . Determining $\Delta_{1}$ in an arbitrary measurement basis is, in general, a difficult task. However, our framework provides a practical way to extract this quantity numerically for any Pauli basis. We apply this method to several measurement bases; in particular, we present more detailed numerical results for the $\sigma^{x}$ basis.
The Hamiltonian of the transverse-field Ising model is given by
$H=-\sum_{i=1}^{L}\left(J\sigma_{i}^{x}\sigma_{i+1}^{x}+h\sigma_{i}^{z}\right),$ (19)
where $L$ is the system size and the critical point occurs at $h=J$. The corresponding $\mathbf{R}$-matrix for the critical Ising chain is presented in TKR2024 . We now examine the distance-dependent decay of $\mathcal{E}_{A_{1}}^{(\alpha,\mathcal{S}_{B})}$, as defined in Eqs. (14)–(16).
Figure 2 displays the behavior of $\mathcal{E}_{A_{1}}^{(\alpha,\mathcal{S}_{B})}(d)$ for the measurement outcomes $|B_{1}\rangle=|+\rangle^{\otimes L_{B_{1}}}$ and $|B_{2}\rangle=|+\rangle^{\otimes L_{B_{2}}}$, in both the $\sigma^{z}$ and $\sigma^{x}$ bases. In both cases, the entanglement entropy exhibits a clear power-law decay with distance, in accordance with Eq. (17). The extracted scaling dimensions are $\Delta_{1}=1/2$ for the $\sigma^{z}$ basis and $\Delta_{1}=2$ for the $\sigma^{x}$ basis.
A broader set of measurement outcomes, along with their corresponding decay exponents $\eta(\alpha)$, is summarized in Table 1. In what follows we always take $\phi\in[0,\pi/2)$; within this range all of the decay exponents we obtain are independent of $\phi$. For all configurations except configuration 4, any measurement basis with $\theta\in(0,\pi/2]$ yields the same decay exponent $\Delta_{1}$ as the special case $\theta=\pi/2$ (the $\sigma^{x}$ basis). By contrast, in configuration 4 one finds $\Delta_{1}=2$ for all $\theta\in(0,\pi/2)$. We suggest that similar exponents can be obtained for a broader range of crystal configurations. For instance, if the base of a crystal configuration has a definite magnetization, we anticipate it will behave similarly to the all-positive or all-negative configurations. Conversely, when the base has zero net magnetization, we expect its behavior to resemble that of the alternating (+)(+-)( + - ) base configuration.
These results underscore the versatility of our approach in characterizing post-measurement entanglement in critical spin chains. Although our primary focus has been on critical spin chains to facilitate comparison with conformal field theory predictions, our approach can readily be applied away from criticality. In such non-critical regimes, one observes an exponential decay of entanglement with increasing distance.
Measurement Output $B_{1}$ Measurement Output $B_{2}$ Scaling dimension $\boldsymbol{(\Delta_{1})}$
1 $|+\rangle_{z}^{\otimes L_{B_{1}}}$ $|+\rangle_{z}^{\otimes L_{B_{2}}}$ 1/2
2 $|+\rangle_{x}^{\otimes L_{B_{1}}}$ $|+\rangle_{x}^{\otimes L_{B_{2}}}$ 2
3 $|+\rangle_{x}^{\otimes L_{B_{1}}}$ $|-\rangle_{x}^{\otimes L_{B_{2}}}$ 1
4 $|\!+\!-\rangle_{x}^{\otimes\frac{L_{B_{1}}}{2}}$ $|\!+\!-\rangle_{x}^{\otimes\frac{L_{B_{2}}}{2}}$ 1/2
5 $|\!+\!-\rangle_{x}^{\otimes\frac{L_{B_{1}}}{2}}$ $|+\rangle_{x}^{\otimes L_{B_{2}}}$ 1
Table 1: Scaling dimensions $\Delta_{1}$ corresponding to different measurement outcome configurations on subsystems $B_{1}$ and $B_{2}$.
VI Conclusion
In this work, we derived an explicit Pfaffian formula for computing the amplitudes of fermionic Gaussian pure states in any Pauli basis, along with a recursive relation that connects amplitudes across different qubit sizes. Together, these results establish a robust computational framework that will be valuable for addressing complex quantum information problems, including formation probabilities, global entanglement, and quantum tomography. For illustration, we calculated the post-measurement entanglement entropy in transverse field Ising chain, uncovering scaling behavior consistent with conformal field theory. Our numerical results yield the precise decay exponents, demonstrating a degree of universality.
Although our analysis exclusively focuses on pure fermionic Gaussian states, it is important to note that many NISQ-era experiments operate under regimes where noise levels are sufficiently low—either naturally or via error mitigation—so that effective purity is maintained over the relevant timescales. In this context, fermionic Gaussian states serve not only as a theoretically tractable subclass of quantum states but also as practically relevant benchmarks for assessing quantum hardware performance Gluza2018 ; Gorman2022 ; Mele2025 . Future work will consider extending these techniques to include mixed-state dynamics and explicit noise models.
@@ VI Conclusion
Applying these formulas to specific quantum systems is a natural next step, expected to advance both theoretical and experimental research in Gaussian state analysis. Open questions remain, especially regarding extensions to Gaussian mixed states, where finding similar explicit formulas would broaden the applicability of Gaussian states in many-body physics and quantum computation. The methods introduced here offer essential tools for enhanced state analysis and practical applications in these fields.
@@
Acknowledgements. We thank D. Brod for discussion. MAR thanks CNPq and FAPERJ (Grant No. E-26/210.062/2023) for partial support. The authors gratefully acknowledge the resources on the LiCCA HPC cluster of the University of Augsburg. The LiCCA cluster is co-funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 499211671.